# Vojta’s abc conjecture for entire curves in toric varieties highly ramified over the boundary

### v1

@@ 5. Proof of the main theorem
Let $D:=\sum_{i=0}^{n}H_{i}$, where $H_{i}:=[x_{i}=0]$, $0\leq i\leq n$, are the coordinate hyperplanes. We consider ${\mathcal{Y}_{f}}:={\mathbb{C}}\backslash f^{-1}(D)$. By (ii) of Proposition 8,
(5.1) $\mathfrak{X}_{\sigma}^{+}(r)\leq\sum_{i=0}^{n}N_{f}^{(1)}(H_{i},r)+O(\log r).$
Since ${\rm mult}_{z_{0}}(f^{*}H_{i})\geq\ell$ for all $i$ with $0\leq i\leq n$ and for every $z_{0}\in{\mathbb{C}}$, we have $N_{f}^{(1)}(H_{i},r)\leq{1\over\ell}T_{f}(r)$ for $i=0,1,\dots,n$. Thus
@@
where $\ell$ is a positive integer to be determined. Let $\tilde{f}_{i}$ be the restriction of $f_{i}$ to $\mathcal{Y}_{f}$ and let $u_{i}:=\frac{\tilde{f}_{i}}{\tilde{f}_{0}}$. Then $u_{i},1\leq i\leq n$, are units on $\mathcal{Y}_{f}$. Denote by ${\tilde{u}}=(1,u_{1},\ldots,u_{n})$ and $\mathbf{u}=(u_{1},\ldots,u_{n})$. Then $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}=T_{\tilde{u},\mathcal{Y}_{f}}(r)+O(1)=T_{{\tilde{f}},\mathcal{Y}_{f}}(r)+O(1),$ where $\tilde{f}$ is the restriction of $f$ to $\mathcal{Y}_{f}$.
We verify that
(5.3) $\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}\geq c_{0}(\mathfrak{X}_{\sigma}^{+}(r)+\log r)$
holds, where $c_{0}$ is a sufficient large constant determined in Theorem 22. Indeed, if it is not true, then there exists a constant $c>0$ and a sequence $r_{n}\rightarrow+\infty$ such that
$\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r_{n})\}\leq c(\mathfrak{X}_{\sigma}^{+}(r_{n})+\log r_{n}).$
Using (ii) of Proposition 8, (5.2) and the fact that $\log r=o(T_{f}(r))$, we have that
$T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})\leq{c(n+2)\over\ell}T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})+o(T_{\tilde{u},\mathcal{Y}_{f}}(r_{n})),$
which gives a contradiction when $\ell>c(n+2)$. Hence (5.3) holds. Therefore, by Theorem 22, there exist a proper Zariski closed subset $W$ of $\mathbb{A}^{n}$ and a positive constant $c$, which are independent of $f$, such that the following two statements hold
(5.4) $N_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)-N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))\leq_{\operatorname{exc}}\frac{\epsilon}{3}\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}.$
(5.5) $N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))\geq_{\operatorname{exc}}\left(\deg G-\frac{\epsilon}{2}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}.$
Here, $\tilde{G}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ such that $\tilde{G}(x_{1},\ldots,x_{n})=G(1,x_{1},\ldots,x_{n})$. Therefore, $\tilde{G}({\mathbf{u}})=G(\tilde{u})$. Since $\widetilde{f}$ is the restriction of $f$ to $\mathcal{Y}_{f}=\mathbb{C}\setminus{f^{-1}(D)}$ and $\tilde{f}_{i}$ has no zero on $\mathcal{Y}_{f}$, it is obvious that
(5.6) $N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)=N^{(1)}_{G(\widetilde{f}),\mathcal{Y}_{f}}(0,r)\leq N^{(1)}_{G(f)}(0,r).$
Then it follows immediately from (5.5) that
$N^{(1)}_{G(f)}(0,r)\geq_{exc}\left(\deg G-\frac{\epsilon}{2}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\}\geq(\deg G-\epsilon)T_{f}(r)$
by (ii) of Proposition 8 and the fact that $\log r=o(T_{f}(r))$. This proves (ii).
To show (i), we apply (5.6) again and obtain the following:
$N_{G(f)}(0,r)-N^{(1)}_{G(f)}(0,r)$ $\leq N_{G(f)}(0,r)-N^{(1)}_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r))$
$=N_{G(f)}(0,r)-N_{\tilde{G}(\mathbf{u}),\mathcal{Y}_{f}}(0,r)+N_{\tilde{G}({\mathbf{u}}),\mathcal{Y}_{f}}(0,r)-N^{(1)}_{\tilde{G}(\mathbf{u}),\mathcal{Y}_{f}}(0,r))$
$\leq_{\operatorname{exc}}\deg G\cdot T_{f}(r)-\left(\deg G-\frac{5\epsilon}{6}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i},\mathcal{Y}_{f}}(r)\},$
by the First Main Theorem, Theorem 6.4, (5.4) and (5.5). Again as $\log r=o(T_{f}(r))$, it yieids
(5.7) $N_{G(f)}(0,r)-N^{(1)}_{G(f)}(0,r)\leq\epsilon T_{f}(r)$
by (i) of Proposition 8. Finally, we can conclude the proof by taking $Z$ to be the Zariski closure of $W$ in $\mathbb{P}^{n}({\mathbb{C}})$. ∎
@@ 6. Toric Varieties
We recall the following setup of finding a natural finite open covering of $X$ from the proof of [13, Theorem 4.4]. Let $\Sigma$ be the fan corresponding to the smooth projective toric variety $X$. Then there is a finite affine covering $\{X_{\eta}\}$ of $X$, where $\eta\in\Sigma$ is a n-dimensional smooth cone with an isomorphism $i_{\eta}:X_{\eta}\to\mathbb{A}^{n}$. This isomorphism restricts to an automorphism of $\mathbb{G}_{m}^{n}$, where we identify $\mathbb{G}_{m}^{n}\subset X_{\eta}$ naturally as a subset of $X$ and $\mathbb{G}_{m}^{n}\subset\mathbb{A}^{n}$ in the standard way such that $\mathbb{A}^{n}\setminus\mathbb{G}_{m}^{n}$ consists of the affine coordinate hyperplanes $\{x_{i}=0\}$, $1\leq i\leq n$. Under this identification, an entire orbifold curve $f$ into $(X,\Delta)$ corresponds to a collection of $n$-tuples of meromorphic functions ${i}_{\eta}(f)=(u_{\eta,1},\ldots,u_{\eta,n})$ such that the zero multiplicity of each $u_{\eta,i}$, $1\leq i\leq n$, is at least $\ell$ if not zero. If $z_{0}\in\mathbb{C}$ is a pole of some $u_{\eta,i}$, $1\leq i\leq n$, then $f(z_{0})$ is in a component of $D_{0}$, the boundary of $X$, with multiplicity
$-\mathop{\mathrm{min}}\{0,v_{z_{0}}(u_{\eta,1}),\ldots,v_{z_{0}}(u_{\eta,n})\}\geq\ell.$
Let $\widetilde{{i}_{\eta}}:X_{\eta}\to\mathbb{P}^{n}({\mathbb{C}})$ be the composition of $i_{\eta}$ with the natural embedding of $\mathbb{A}^{n}$ to $\mathbb{P}^{n}({\mathbb{C}})$. Then $\widetilde{{i}_{\eta}}(f)=(f_{\eta,0},\ldots,f_{\eta,n})=[1:u_{\eta,1}:\cdots:u_{\eta,n}]:\mathbb{C}\to\mathbb{P}^{n}({\mathbb{C}})$, where $f_{\eta,0},\ldots,f_{\eta,n}$ are entire functions with no common zero. Furthermore, the pullback $(\widetilde{{i}_{\eta}}^{-1})^{*}(D|_{X_{\eta}})$ of $D$ to $\mathbb{P}^{n}({\mathbb{C}})$ is defined by some nonconstant homogeneous polynomial $F_{\eta}\in\mathbb{C}[x_{0},\ldots,x_{n}]$ with no monomial and repeated factors and $\deg_{X_{i}}F_{\eta}=\deg F_{\eta}$ for each $0\leq i\leq n$ since $D$ is reduced and is in general position with the boundary of $\mathbb{G}_{m}^{n}$ in $X$. By [25, Proposition 12.11], there exist a non-zero constant $c_{\eta,A}$ and a proper closed subset $Z_{\eta,A}\subset X_{\eta}$, depending on $\eta$ and $A$ such that
@@
Since $R$ is a big divisor on $Y$, by [25, Proposition 12.11], there exists a constant $c$ and a proper Zariski-closed subset $W_{0}$ of $Y$, depending only on $A$ and $R$, such that such that

### v2

@@ 5. Proof of the main theorem
Let $D:=\sum_{i=0}^{n}H_{i}$, where $H_{i}:=[x_{i}=0]$, $0\leq i\leq n$, are the coordinate hyperplanes. We consider ${\mathcal{Y}_{f}}:={\mathbb{C}}\backslash f^{-1}(D)$ with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$, as being constructed in Section 3.1. Furthermore, by further shrinking $r_{j}$, we may assume that
(5.1) $|T_{u_{i}}(r)-T^{M}_{u_{i},\sigma}(r)|\leq\log r+O(1)$
for $1\leq i\leq n$ by (3.15) and Proposition 12 (ii), and
(5.2) $\mathfrak{X}_{\sigma}^{+}(r)\leq\sum_{i=0}^{n}N_{f}^{(1)}(H_{i},r)+O(\log r)$
by (3.17). Since ${\rm mult}_{z_{0}}(f^{*}H_{i})\geq\ell$ for all $i$ with $0\leq i\leq n$ and for every $z_{0}\in{\mathbb{C}}$, we have $N_{f}^{(1)}(H_{i},r)\leq{1\over\ell}T_{f}(r)$ for $i=0,1,\dots,n$. Thus
@@
where $\ell$ is a positive integer to be determined.
Let $u_{i}:=\frac{f_{i}}{f_{0}}$. Then $u_{i},1\leq i\leq n$, are units on $\mathcal{Y}_{f}$. Let $\tilde{G}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ be defined by $\tilde{G}(x_{1},\ldots,x_{n})=G(1,x_{1},\ldots,x_{n})$. Then $f_{0}^{\deg G}\cdot{\tilde{G}}({\mathbf{u}})=G(f).$ Hence it is obvious that
(5.4) $N^{(1)}_{\tilde{G}({\mathbf{u}}),\sigma}(0,r)=N^{(1)}_{G(f),\sigma}(0,r).$
Let $\epsilon>0$, and let $W$ be the proper Zariski closed subset $W$ of $\mathbb{A}^{n}$ determined in Theorem 26. We express the zero locus of $W$ as the vanishing set of a finite collection $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$, and assume that $H({\bf u})\neq 0$ for all $H\in\Sigma$. Then, by Theorem 26, together with (5.1) and (5.4), we obtain, after shrinking $r_{j}$ if necessary, that there exists a positive real $c_{0}$ (depending only on $\epsilon$, $n$, and $\deg G$) such that
(5.5) $N_{G(f),\sigma}(0,r)-N^{(1)}_{G(f),\sigma}(0,r)\leq_{\operatorname{exc}}\frac{\epsilon}{3}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}+c_{0}(\mathfrak{X}_{\sigma}^{+}(r)+\log r),$
and
(5.6) $\left(\deg G-\frac{\epsilon}{2}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}\leq_{\operatorname{exc}}N^{(1)}_{G(f),\sigma}(0,r)+c_{0}(\mathfrak{X}_{\sigma}^{+}(r)+\log r).$
Using (5.3) and the fact that $\log r=o(T_{f}(r))$, by taking $\ell>3c_{0}(n+2)\epsilon^{-1}$, we have that
(5.7) $c_{0}(\mathfrak{X}_{\sigma}^{+}(r)+\log r)\leq{\epsilon/3}\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}.$
Since $B_{t}^{\sigma}\subset\mathbb{D}_{t}$ for any $t\geq 1$, $N^{(1)}_{G(f),\sigma}(0,r)\leq N^{(1)}_{G(f)}(0,r)$. It follows immediately from (5.6) and (5.7) that
(5.8) $N^{(1)}_{G(f)}(0,r)\geq_{\rm exc}\left(\deg G-\frac{5\epsilon}{6}\right)\cdot\max_{1\leq i\leq n}\{T_{u_{i}}(r)\}\geq(\deg G-\epsilon)T_{f}(r).$
This proves (ii).
It remains to prove (i), which follows from (ii), i.e., from (5.8) together with the First Main Theorem.
$N_{G(f)}(0,r)-N^{(1)}_{G(f)}(0,r)$ $\leq_{\operatorname{exc}}\deg G\cdot T_{f}(r)-(\deg G-\epsilon)T_{f}(r)\leq_{\operatorname{exc}}\epsilon T_{f}(r).$
Finally, we can conclude the proof by taking $Z$ to be the Zariski closure of $W$ in $\mathbb{P}^{n}({\mathbb{C}})$. ∎
@@ 6. Toric Varieties
We recall the following setup of finding a natural finite open covering of $X$ from the proof of [12, Theorem 4.4]. Let $\Sigma$ be the fan corresponding to the smooth projective toric variety $X$. Then there is a finite affine covering $\{X_{\eta}\}$ of $X$, where $\eta\in\Sigma$ is a n-dimensional smooth cone with an isomorphism $i_{\eta}:X_{\eta}\to\mathbb{A}^{n}$. This isomorphism restricts to an automorphism of $\mathbb{G}_{m}^{n}$, where we identify $\mathbb{G}_{m}^{n}\subset X_{\eta}$ naturally as a subset of $X$ and $\mathbb{G}_{m}^{n}\subset\mathbb{A}^{n}$ in the standard way such that $\mathbb{A}^{n}\setminus\mathbb{G}_{m}^{n}$ consists of the affine coordinate hyperplanes $\{x_{i}=0\}$, $1\leq i\leq n$. Under this identification, an entire orbifold curve $f$ into $(X,\Delta)$ corresponds to a collection of $n$-tuples of meromorphic functions ${i}_{\eta}(f)=(u_{\eta,1},\ldots,u_{\eta,n})$ such that the zero multiplicity of each $u_{\eta,i}$, $1\leq i\leq n$, is at least $\ell$ if not zero. If $z_{0}\in\mathbb{C}$ is a pole of some $u_{\eta,i}$, $1\leq i\leq n$, then $f(z_{0})$ is in a component of $D_{0}$, the boundary of $X$, with multiplicity
$\mathop{\mathrm{min}}\{0,v_{z_{0}}(u_{\eta,1}),\ldots,v_{z_{0}}(u_{\eta,n})\}\geq\ell.$
Let $\widetilde{{i}_{\eta}}:X_{\eta}\to\mathbb{P}^{n}({\mathbb{C}})$ be the composition of $i_{\eta}$ with the natural embedding of $\mathbb{A}^{n}$ to $\mathbb{P}^{n}({\mathbb{C}})$. Then $\widetilde{{i}_{\eta}}(f)=(f_{\eta,0},\ldots,f_{\eta,n})=[1:u_{\eta,1}:\cdots:u_{\eta,n}]:\mathbb{C}\to\mathbb{P}^{n}({\mathbb{C}})$, where $f_{\eta,0},\ldots,f_{\eta,n}$ are entire functions with no common zero. Furthermore, the pullback $(\widetilde{{i}_{\eta}}^{-1})^{*}(D|_{X_{\eta}})$ of $D$ to $\mathbb{P}^{n}({\mathbb{C}})$ is defined by some nonconstant homogeneous polynomial $F_{\eta}\in\mathbb{C}[x_{0},\ldots,x_{n}]$ with no monomial and repeated factors and $\deg_{X_{i}}F_{\eta}=\deg F_{\eta}$ for each $0\leq i\leq n$ since $D$ is reduced and is in general position with the boundary of $\mathbb{G}_{m}^{n}$ in $X$. By [27, Proposition 12.11], there exist a non-zero constant $c_{\eta,A}$ and a proper closed subset $Z_{\eta,A}\subset X_{\eta}$, depending on $\eta$ and $A$ such that
@@
Since $R$ is a big divisor on $Y$, by [27, Proposition 12.11], there exists a constant $c$ and a proper Zariski-closed subset $W_{0}$ of $Y$, depending only on $A$ and $R$, such that such that
@@
Acknowledgements. We sincerely thank the referees for their careful reading of the manuscript and for their many valuable comments and suggestions. We also thank the first author’s student, Zhe Wang, for helpful discussions and valuable suggestions.